OpenStudy (anonymous):

Please help Please Part 1: Solve the inequality 2(x + 6) greater than or equal to 3x - 2 and show all of your work. You can use >= to represent the greater than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (anonymous):

2(x+6) >= 3x-2 do you know how to start to solve for x?

OpenStudy (anonymous):

first distribute the 2 so 2(x+6) >= 3x-2 2x+12>=3x-2 then add 2 to both sides i thnk so 2x+14>=3 so no subtract the 2 i think so 14>=x

OpenStudy (anonymous):

2(x+6)>=3x-2 2x+12>=3x-2 3x-2x<=12+2 x<=14 close circle on 14:shading to the left

OpenStudy (anonymous):

Thank you =) so much i had it all wrong

OpenStudy (anonymous):

but you DID get it right!